# Disambiguation for Video Frame Interpolation

### v3

@@ Abstract
Index Terms:
Video frame interpolation, Temporal super-resolution, Disambiguation, Video editing
@@ I Introduction
Although distance indexing addresses the scalar speed ambiguity, the directional ambiguity of motion remains a challenge. We observed that this directional uncertainty is most pronounced when situated equally far from the two input frames, i.e., halfway between them. Drawing inspiration from countless computer vision algorithms that iteratively solve a difficult problem (e.g., optical flow [9] and image generation [10]), we introduce an iterative reference-based estimation strategy. This strategy seeks to mitigate directional ambiguity by incrementally estimating distances, beginning with nearby points and advancing to farther ones, such that the uncertainty within each step is minimized and the image quality is further improved.
@@ II-A Video frame interpolation
II-A1 General overview.
Numerous VFI solutions rely on optical flows to predict latent frames. Typically, these methods warp input frames forward or backward using flow calculated by off-the-shelf networks like  [14, 15, 16, 9] or self-contained flow estimators like [7, 17, 18]. Networks then refine the warped frame to improve visual quality. SuperSlomo [6] uses a linear combination of bi-directional flows for intermediate flow estimation and backward warping. DAIN [1] introduces a depth-aware flow projection layer for advanced intermediate flow estimation. AdaCoF [19] estimates kernel weights and offset vectors for each target pixel, while BMBC [20] and ABME [21] refine optical flow estimation. Large motion interpolation is addressed by XVFI [22] through a recursive multi-scale structure. VFIFormer [23] employs Transformers to model long-range pixel correlations. IFRNet [24], RIFE [7], and UPR-Net [25] employ efficient pyramid network designs for high-quality, real-time interpolation, with IFRNet and RIFE using leakage distillation losses for flow estimation. Recently, more advanced network modules and operations are proposed to push the upper limit of VFI performance, such as the transformer-based bilateral motion estimator of BiFormer [26], a unifying operation of EMA-VFI [17] to explicitly disentangle motion and appearance information, and bi-directional correlation volumes for all pairs of pixels of AMT [18]. On the other hand, SoftSplat [27] and M2M [28] actively explore the forward warping operation for VFI. More recently, diffusion-based video frame interpolators, such as LDMVFI [29] and SVDKFI [30], have leveraged generative priors to enhance visual quality but still face challenges such as high computational costs and slow runtimes.
Other contributions to VFI come from various perspectives. For instance, Xu et al. [8, 31] leverage acceleration information from nearby frames, VideoINR [32] is the first to employ an implicit neural representation, and Lee et al. [33] explore and address discontinuity in video frame interpolation using figure-text mixing data augmentation and a discontinuity map. Flow-free approaches have also attracted interest. SepConv [34] integrates motion estimation and pixel synthesis, CAIN [35] employs the PixelShuffle operation with channel attention, and FLAVR [36] utilizes 3D space-time convolutions. Additionally, specialized interpolation methods for anime, which often exhibit minimal textures and exaggerated motion, are proposed by AnimeInterp [37] and Chen et al. [38]. On the other hand, motion induced blur [39, 40, 41], shutter mode [42, 43, 44], and event camera [45, 46] are also exploited to achieve VFI.
@@ III Velocity Ambiguity
Although the starting and ending positions of the baseball are given, its location at $t=0.5$ remains ambiguous due to an unknown speed distribution: The ball can be accelerating or decelerating, resulting in different locations. This ambiguity introduces a challenge in model training as it leads to multiple valid supervision targets for the identical input. Contrary to the deterministic scenario illustrated in Eq. 1, the VFI function $\mathcal{F}$ is actually tasked with generating a distribution of plausible frames from the same input frames and time indexing. This can be expressed as:
@@
[…]
@@ IV-A Distance indexing
$D_{t}(x,y)=\frac{\left\|\mathbf{V}_{0\to t}(x,y)\right\|\cos{\theta}}{\left\|\mathbf{V}_{0\to 1}(x,y)\right\|},$ (7)
@@ IV-B Iterative reference-based estimation
where $i\in\{0,1,\dots,N-1\}$. For example, if we break the estimation of of a remote step $t$ into two steps:
@@ V-B Multi-frame refiner
${I^{\prime}_{t}=M_{1}\odot I_{t}^{{}^{\prime}+}+M_{2}\odot I_{t}+M_{3}\odot I_{t}^{{}^{\prime}-}}.$ (18)
@@
Figure 6: Comparison of qualitative results. $[T]$: original arbitrary time VFI models using time indexing. $[D]_{u}$: models trained using our distance indexing, then inference using uniform maps. $[T,R]$: models using time indexing with iterative reference-based estimation. $[D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps.
@@ V-C Model tuning
During training, we first freeze the parameters of the two-frame VFI model $\mathcal{F}$ and the continuous map estimator. Only the refiner $\mathcal{F^{\prime}}$ is updated during this stage, using the same training configuration as the original model. This setup enables the refiner to learn how to enhance interpolation by leveraging information from more distant adjacent frames. Next, we freeze the parameters of the pretrained map estimator and jointly optimize both the VFI model $\mathcal{F}$ and the proposed refiner $\mathcal{F^{\prime}}$. This approach enables the entire system to adapt to indexing maps derived from optical flow predicted by CPFlow. We experimented with various training strategies and loss functions to evaluate their effectiveness (see Section VI-H).
@@ VI-A Implementation
We leveraged the plug-and-play nature of our distance indexing and iterative reference-based estimation strategies to seamlessly integrate them into influential arbitrary time VFI models such as RIFE [7] and IFRNet [24], and state-of-the-art models including AMT [18] and EMA-VFI [17]. We adhere to the original hyperparameters for each model for a fair comparison and implement them with PyTorch [53]. For training, we use the septuplet dataset from Vimeo90K [54]. The septuplet dataset comprises 91,701 seven-frame sequences at $448\times 256$, extracted from 39,000 video clips. For evaluation, we use both pixel-centric metrics like PSNR and SSIM [55], and perceptual metrics such as reference-based LPIPS [56] and non-reference NIQE [57]. Concerning the iterative reference-based estimation strategy, $D_{ref}$ during training is calculated from the optical flow derived from ground-truth data at a time point corresponding to a randomly selected reference frame, like $t/2$. In the inference phase, we similarly employ a uniform map for reference, for example, setting $D_{ref}=t/2$.
@@ VI-B Qualitative comparison
VI-C Quantitative comparison
VI-C1 Convergence curves
VI-C2 Comparison on Vimeo90K septuplet dataset
TABLE I: Comparison on Vimeo90K septuplet dataset. $[T]$ denotes the method trained with traditional arbitrary time indexing paradigm. $[D]$ and $[R]$ denote the distance indexing paradigm and iterative reference-based estimation strategy, respectively. $[R]$ uses 2 iterations by default. $[\cdot]_{u}$ denotes inference with uniform map as time indexes. We utilize the first and last frames as inputs to predict the rest five frames. The
@@ VI-C2 Comparison on Vimeo90K septuplet dataset
TABLE II: Ablation study of the number of iterations on Vimeo90K septuplet dataset. $[\cdot]^{\#}$ denotes the number of iterations used for inference.
@@
$[D,R]_{u}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$
@@
$[T,R]$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$ $[\cdot]^{1}$ $[\cdot]^{2}$ $[\cdot]^{3}$
@@ VI-C3 Ablation study of the number of iterations
Table II offers an ablation study on the number of iterations and the efficacy of a pure iterative reference-based estimation strategy. The upper section suggests that setting iterations at two strikes a good trade-off between computational efficiency and performance. The lower segment illustrates that while iterative reference-based estimation generally works for time indexing, there are exceptions, as observed with AMT-S. However, when combined with distance indexing, iterative reference-based estimation exhibits more stable improvement, as evidenced by the results for $[D,R]_{u}$. This is consistent with qualitative comparison.
TABLE III: Comparison on Adobe240 [58] for $\times$8 interpolation with RIFE [7].
$[T]$ $[D]_{u}$ $[D,R]_{u}$
PSNR $\uparrow$ 30.24
30.47
30.30
SSIM $\uparrow$
0.939
0.938 0.937
LPIPS $\downarrow$ 0.073 0.057
0.054
NIQE $\downarrow$ 5.206 4.974
4.907
TABLE IV: Comparison on X4K1000FPS [22] for $\times$8 interpolation with RIFE [7].
$[T]$ $[D]_{u}$ $[D,R]_{u}$
PSNR $\uparrow$ 36.36
36.80
36.26
SSIM $\uparrow$
0.967
0.964 0.964
LPIPS $\downarrow$ 0.040 0.032
0.032
NIQE $\downarrow$ 7.130 6.936
6.924
TABLE V: Comparison on X4K1000FPS [22] for $\times$16 interpolation with RIFE [7].
$[T]$ $[D]_{u}$ $[D,R]_{u}$
@@
TABLE VI: Comparison on Vimeo90K [54] using GMFlow [59] for distance map calculation with RIFE [7].
$[T]$ $[D]_{u}$ $[D,R]_{u}$
@@
VI-C4 Comparison on other benchmarks
The septuplet set of Vimeo90K [54] is large enough to train a practical video frame interpolation model, and it represents the situations where the temporal distance between input frames is large. Thus, Vimeo90K (septuplet) can well demonstrate the velocity ambiguity problem that our work aims to highlight. Nonetheless, we report more results on other benchmarks. Table VI and Table VI show the results of RIFE [7] on Adobe240 [58] and X4K1000FPS [22] for $\times$8 interpolation respectively, using uniform maps. Distance indexing $[D]_{u}$ and iterative reference-based estimation $[R]_{u}$ strategies can consistently help improve the perceptual quality. In addition, it is noteworthy that $[D]_{u}$ is better than $[T]$ in terms of the pixel-centric metrics like PSNR, showing that the constant speed assumption (uniform distance maps) holds well on these two easier benchmarks. We further show $\times$16 interpolation on X4K1000FPS with larger temporal distance in Table VI. The results highlight that the benefits of our strategies are more pronounced with increased temporal distances.
@@ VI-C5 Other optical flow estimator
We also employ GMFlow [59] for the precomputation of distance maps, enabling an analysis of model performance when integrated with alternative optical flow estimations. The results are as shown in Table VI. Our strategies still lead to consistent improvement on perceptual metrics. However, this more recent and performant optical flow estimator does not introduce improvement compared to RAFT [9]. A likely explanation is that since we quantify the optical flow to $[0,1]$ scalar values for better generalization, our training strategies are less sensitive to the precision of the optical flow estimator.
VI-D User study
VI-E 2D manipulation of frame interpolation

### v4

# Velocity Disambiguation for Video Frame Interpolation

@@ I Introduction
Although distance indexing addresses the scalar speed ambiguity, the directional ambiguity of motion remains a challenge. Empirically we found that, although learning-based video frame interpolation can handle minor directional uncertainty at a small timestep, the ambiguity becomes pronounced at the temporal midway between the two input frames, i.e., t=0.5, as illustrated in Fig. 3(b). Inspired by iterative inference paradigms in optical flow [49] and image generation [40], we introduce an iterative reference-based estimation strategy. Rather than estimating the full motion field at once, our approach decomposes the problem into incremental distance steps. By propagating estimates from nearby to farther points, we constrain the search space at each iteration, thereby minimizing directional uncertainty and enhancing synthesis quality.
@@ II-A Video frame interpolation
II-A1 General overview
Numerous VFI solutions rely on optical flows to predict latent frames. Typically, these methods warp input frames forward or backward using flow calculated by off-the-shelf networks like  [48, 8, 16, 49] or self-contained flow estimators like [14, 65, 26]. Networks then refine the warped frame to improve visual quality. SuperSlomo [18] uses a linear combination of bi-directional flows for intermediate flow estimation and backward warping. DAIN [1] introduces a depth-aware flow projection layer for advanced intermediate flow estimation. AdaCoF [24] estimates kernel weights and offset vectors for each target pixel, while BMBC [35] and ABME [36] refine optical flow estimation. Large motion interpolation is addressed by XVFI [44] through a recursive multi-scale structure. VFIFormer [29] employs Transformer architectures to model long-range pixel correlations, while VFIMamba [64] adopts a Mamba-based architecture for efficient sequence modeling. IFRNet [22], RIFE [14], and UPR-Net [19] employ efficient pyramid network designs for high-quality, real-time interpolation, with IFRNet and RIFE using leakage distillation losses for flow estimation. Recently, more advanced network modules and operations are proposed to push the upper limit of VFI performance, such as the transformer-based bilateral motion estimator of BiFormer [34], a unifying operation of EMA-VFI [65] to explicitly disentangle motion and appearance information, and bi-directional correlation volumes for all pairs of pixels of AMT [26]. On the other hand, SoftSplat [32] and M2M [13] actively explore the forward warping operation for VFI. To further improve perceptual clarity in video frame interpolation, FILM [39] introduces perceptual losses during training, while uncertainty-aware and adaptive interpolation methods have also been proposed to better handle ambiguous regions [38, 2]. More recently, diffusion-based video frame interpolators, such as LDMVFI [7] and SVDKFI [53], have leveraged generative priors to enhance visual quality but still face challenges such as high computational costs and slow runtimes.
Other contributions to VFI come from various perspectives. For instance, Xu et al. [59, 12] leverage acceleration information from nearby frames, VideoINR [4] is the first to employ an implicit neural representation, and Lee et al. [25] explore and address discontinuity in video frame interpolation using figure-text mixing data augmentation and a discontinuity map. Flow-free approaches have also attracted interest. SepConv [33] integrates motion estimation and pixel synthesis, CAIN [6] employs the PixelShuffle operation with channel attention, and FLAVR [20] utilizes 3D space-time convolutions. Additionally, specialized interpolation methods for anime, which often exhibit minimal textures and exaggerated motion, are proposed by AnimeInterp [46] and Chen et al. [3]. On the other hand, motion induced blur [42, 70, 67], shutter mode [9, 68, 17], and event camera [50, 27] are also exploited to achieve VFI. For a more comprehensive overview of recent advances in video frame interpolation, we refer readers to the survey by Kye et al. [23].
@@ III Velocity Ambiguity
Although the starting and ending positions of the baseball are given, its location at $t=0.5$ remains ambiguous due to an unknown speed distribution: The ball can be accelerating or decelerating, resulting in different locations. This ambiguity introduces a challenge in model training as it leads to multiple valid supervision targets for the identical input. Contrary to the deterministic scenario illustrated in Eq. 1, the VFI function $\mathcal{F}$ is actually tasked with generating a sampled sequence of plausible frames within the distribution derived from the same input frames and time indexing. This can be expressed as:
@@
$\hat{I}_{t}=\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[I_{t}].\qquad\text{(See details in Appendix)}$ (4)
@@ IV-A Distance indexing
$D_{t}(x,y)=\frac{\mathbf{V}_{0\to t}(x,y)\cdot\mathbf{V}_{0\to 1}(x,y)}{\left\|\mathbf{V}_{0\to 1}(x,y)\right\|^{2}},$ (7)
@@ IV-B Iterative reference-based estimation
where $i\in\{0,1,\dots,N-1\}$. For example, if we break the estimation of a remote step $t$ into two steps:
@@ V-B Multi-frame refiner
$I^{\prime}_{t}=M_{1}\odot I_{t}^{{}^{\prime}+}+M_{2}\odot I_{t}+M_{3}\odot I_{t}^{{}^{\prime}-}$} (18)
@@
Figure 6: Comparison of qualitative results. $[T]$: original arbitrary time VFI models using time indexing. $D]_{u}$: models trained using our distance indexing, then inference using uniform maps. $[T,R]$: models using time indexing with iterative reference-based estimation. $D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps.
@@ V-C Model tuning
During training, we first freeze the parameters of the two-frame VFI model $\mathcal{F}$ and the continuous map estimator. Only the refiner $\mathcal{F^{\prime}}$ is updated during this stage, using the same training configuration as the original model. This setup enables the refiner to learn how to enhance interpolation by leveraging information from more distant adjacent frames. Next, we freeze the parameters of the pretrained map estimator and jointly optimize both the VFI model $\mathcal{F}$ and the proposed refiner $\mathcal{F^{\prime}}$. This approach enables the entire system to adapt to indexing maps derived from optical flow predicted by CPFlow. We experimented with various training strategies and loss functions to evaluate their effectiveness (see Section VI-I).
@@ VI-A Implementation
We leveraged the plug-and-play nature of our distance indexing and iterative reference-based estimation strategies to seamlessly integrate them into influential arbitrary time VFI models such as RIFE [14] and IFRNet [22], and state-of-the-art models including AMT [26] and EMA-VFI [65]. To further strengthen the empirical evidence, we additionally validated the effectiveness of distance indexing on a diffusion-based model, LDMVFI [7], and a representative multi-frame method, VFI-Transformer [43]. We adhere to the original hyperparameters for each model for a fair comparison and implement them with PyTorch [37]. For training, we use the septuplet dataset from Vimeo90K [61]. The septuplet dataset comprises 91,701 seven-frame sequences at $448\times 256$, extracted from 39,000 video clips. For evaluation, we use both pixel-centric metrics like PSNR and SSIM [55], and perceptual metrics such as reference-based LPIPS [66] and non-reference NIQE [31]. Concerning the iterative reference-based estimation strategy, $D_{ref}$ during training is calculated from the optical flow derived from ground-truth data at a time point corresponding to a randomly selected reference frame, like $t/2$. In the inference phase, we similarly employ a uniform map for reference, for example, setting $D_{ref}=t/2$.
@@ VI-B2 User study
TABLE I: Comparison on Vimeo90K septuplet dataset. $[T]$ denotes the method trained with traditional arbitrary time indexing paradigm. $[D]$ and $[R]$ denote the distance indexing paradigm and iterative reference-based estimation strategy, respectively. $[R]$ uses 2 iterations by default. $\cdot]_{u}$ denotes inference with uniform map as time indexes. We utilize the first and last frames as inputs to predict the rest five frames. The
@@
VI-C Quantitative comparison
VI-C1 Convergence curves
VI-C2 Comparison on Vimeo90K septuplet dataset
TABLE II: Ablation study of the number of iterations on Vimeo90K septuplet dataset. $\cdot$$^{\#}$ denotes the number of iterations used for inference.
@@ VI-C2 Comparison on Vimeo90K septuplet dataset
$D,R]_{u} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3}$
@@
$[T,R]$ $\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3} [$$\cdot]^{1} [$$\cdot]^{2} [$$\cdot]^{3}$
@@ VI-C3 Ablation study of the number of iterations
Table II offers an ablation study on the number of iterations and the efficacy of a pure iterative reference-based estimation strategy. The upper section suggests that setting iterations at two strikes a good trade-off between computational efficiency and performance. The lower segment illustrates that while iterative reference-based estimation generally works for time indexing, there are exceptions, as observed with AMT-S. However, when combined with distance indexing, iterative reference-based estimation exhibits more stable improvement, as evidenced by the results for $D,R]_{u}$. This is consistent with qualitative comparison.
TABLE III: Comparison on X4K1000FPS [44] for $\times$16 interpolation with RIFE [14].
$[T]$ $D]_{u} [$$D,R]_{u}$
@@
VI-C4 Comparison on other benchmarks
The septuplet set of Vimeo90K [61] is large enough to train a practical video frame interpolation model, and it represents the situations where the temporal distance between input frames is large. Thus, Vimeo90K (septuplet) can well demonstrate the velocity ambiguity problem that our work aims to highlight. We further show $\times$16 interpolation on X4K1000FPS with larger temporal distance in Table III. The results highlight that the benefits of our strategies are more pronounced with increased temporal distances.
TABLE IV: Comparison on Vimeo90K [61] using GMFlow [58] for distance map calculation with RIFE [14].
$[T]$ $D]_{u} [$$D,R]_{u}$
@@ VI-C5 Other optical flow estimator
We also employ GMFlow [58] for the precomputation of distance maps, enabling an analysis of model performance when integrated with alternative optical flow estimations. The results are as shown in Table IV. Our strategies still lead to consistent improvement on perceptual metrics. However, this more recent and performant optical flow estimator does not introduce improvement compared to RAFT [49]. A likely explanation is that since we quantify the optical flow to $[0,1]$ scalar values for better generalization, our training strategies are less sensitive to the precision of the optical flow estimator.
TABLE V: Comparison on the septuplet of Vimeo90K [61] using LPIPS loss [66]. We use RIFE [14] as a representative example.
$[T]$ $D]_{u} [$$D,R]_{u}$
PSNR $\uparrow$
27.19
26.71 26.72
SSIM $\uparrow$
0.898
0.889 0.890
LPIPS $\downarrow$
0.061
0.065 0.064
NIQE $\downarrow$ 6.307 5.901
5.837
VI-C6 Comparison of using perceptual loss
In addition to training with traditional pixel losses based on L1 and L2 losses, we present the results of employing the more recent LPIPS loss [66] with a VGG backbone [45], as shown in Table V. The non-reference perceptual quality metric, NIQE, shows notable improvement across all variants. The results also consistently demonstrate the effectiveness of our strategies in resolving velocity ambiguity. Besides, due to the direct optimization of LPIPS loss, the assumption of constant speed in uniform maps affects the performance for this metric. This is why in the test results, $[T]$ has a lower LPIPS, while $D]_{u}$ and $D,R]_{u}$ are slightly higher.
TABLE VI: Temporal consistency evaluation on Vimeo90K Septuplet using FVD [51] and VBench [15].
RIFE [14] IFRNet [22] AMT-S [26] EMA-VFI [65]
$[T]$ $D]_{u} [$T][T] $D]_{u} [$T][T] $D]_{u} [$T][T] $D]_{u}$
FVD $\downarrow$ 0.0174
0.0137
0.0142
0.0137
0.0167
0.0137
0.0119
0.0108
Subject Consistency $\uparrow$ 0.957
0.959
0.958
0.961
0.958
0.960
0.962
0.967
Background Consistency $\uparrow$ 0.952
0.955
0.952
0.956
0.953
0.955
0.955
0.958
Imaging Quality $\uparrow$ 0.448
0.475
0.460
0.481
0.462
0.485
0.468
0.486
Motion Smoothness $\uparrow$
0.996
0.996
0.996
0.996
0.996
0.996
0.996
0.996
VI-C7 Temporal consistency
We have further evaluated models using FVD [51] and VBench [15]. As shown in Table VI, replacing time indexing with distance indexing consistently improves FVD scores as well as VBench metrics related to subject consistency, background consistency, and overall imaging quality. We notice that the motion smoothness metric in VBench [15] shows limited discriminative power for video frame interpolation. These results indicate that reducing velocity ambiguity not only enhances per-frame perceptual quality but also leads to more temporally coherent and stable video interpolation.
Figure 9: Comparison on Vimeo90K Septuplet with LDMVFI [7].
TABLE VII: Comparison on Vimeo90K Septuplet with LDMVFI [7].
$[T]$ $[D]$ $D]_{u}$
[…]
27.12
25.94
[…]
0.906
0.893
[…]
0.086
0.090
[…]
6.451
VI-D Evaluating distance indexing on diffusion-based baseline
We further evaluate diffusion-based VFI models, which are known for their strong generative capacity. Since SVDKFI [53] directly generates a fixed sequence of multiple frames in a single pass, we adopt LDMVFI [7] for evaluation, whose formulation is more compatible with our interpolation setting. As reported in Table VII, even though diffusion models may partially mitigate ambiguity through generative priors, replacing time indexing with distance indexing still leads to further improvements. We also provide a qualitative comparison in Fig. 9. These results indicate that velocity ambiguity remains relevant for diffusion-based approaches and that distance indexing provides complementary benefits.
Figure 10: Comparison on Vimeo90K Septuplet with VFI-Transformer [65].
TABLE VIII: Comparison on Vimeo90K Septuplet with VFI-Transformer [65].
$[T]$ $[D]$ $D]_{u}$
[…]
36.46
35.43
[…]
0.976
0.964
[…]
0.079
0.081
[…]
6.082
6.197
VI-E Evaluating distance indexing on multi-frame baseline
We also evaluate a representative multi-frame input method, Video Frame Interpolation Transformer (VFI-Transformer) [43]. As shown in Table VIII and Fig. 10, distance indexing also improves performance in the multi-frame setting, confirming that additional temporal context alone does not fully eliminate velocity ambiguity and that the proposed strategy remains beneficial.
VI-F 2D manipulation of frame interpolation